Marvin's age right now is exactly 3 times his age 3 years in the future minus3 times his age 3 years ago. how old is Marvin right now?
Will this situation occur again during Marvin's life?

Answer by ankor@dixie-net.com(22740)

Marvin's age right now is exactly 3 times his age 3 years in the future minus3 times his age 3 years ago.
how old is Marvin right now?
:
let a = his age right now
:
a = 3(a+3) - 3(a-3)
a = 3a + 9 - 3a + 9
a = 9 + 9
a = 18 yr old is his age now.
:
:
See if that checks out
18 = 3(21) - 3(15)
18 = 63 - 45
:
Will this situation occur again during Marvin's life?
No, only at age 18 (the difference will always by 18 whatever the age)